OpenStudy (anonymous):

(sen+cos)`2+(sen-cos)^2=2(tan^2cos^2+cot^2) help me pls

OpenStudy (anonymous):

(sen+cos)`2+(sen-cos)^2=2(tan^2cos^2+cot^2)

OpenStudy (mathstudent55):

Is `2 supposed to be ^2 ?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

\((\sin x + \cos x)^2+(\sin x - \cos x)^2=2(\tan^2 x \cos^2 x+\cot^2 x)\) Are you solving an equation or proving an identity?

OpenStudy (anonymous):

Trigonometric Identities

OpenStudy (mathstudent55):

Let's try.

OpenStudy (anonymous):

pls ty

OpenStudy (mathstudent55):

\((\sin x + \cos x)^2+(\sin x - \cos x)^2=2(\tan^2 x \cos^2 x+\cot^2 x)\) Square both quantities on the left side: \(\sin^2 x + 2\sin x\cos x + \cos^2 x + \sin^2 x - 2 \sin x \cos x + \cos^2 x\)

OpenStudy (mathstudent55):

\(\sin^2 x + \cos^2 x + \sin^2 x + \cos^2 x\)

OpenStudy (mathstudent55):

\(2(\sin^2 x + \cos^2 x) \) \(2(1)\) \(2\)

OpenStudy (mathstudent55):

The left sides simplifies to just 2. Now let's see what we can do to the right side.

OpenStudy (mathstudent55):

Did you copy the problem correctly? Are you sure on the right side it's cot^2 x and not cos^2 x at the end?

OpenStudy (anonymous):

is cot^2x

OpenStudy (mathstudent55):

Then the identity will not work.

OpenStudy (mathstudent55):

Let's work on the right side.

OpenStudy (mathstudent55):

\(2(\tan^2 x \cos^2 x+\cot^2 x)\) \(2(\dfrac{\sin^2 x}{\cos^2 x} \cos^2 x+\cot^2 x)\) \(2(\sin^2 x+\cot^2 x)\)

OpenStudy (mathstudent55):

Notice that if you had \(\cos ^2 x\) instead of \(\cot^2 x\), then \(\sin^2 x + \cos^2 x\) would add up to 1, and you'd end up with 2 on the right side, but if the right side really has \(\cot^2 x\) and not \(\cos^2 x\), then this is not an identity.

OpenStudy (anonymous):

:/ i dont know in my exercise is cot^2x in the end

OpenStudy (mathstudent55):

Ok. Then it's not an identity. We proved that it is not an identity.